Grade 12Mechanics

A player catches a ball of 200g moving with a speed of 20m/s. If the time taken to complete the catch is 0.5 sec, the force exerted on the players hand is-
A) 8N
B) 4N
C) 2N
D) 0N

Kaviya
7 Years agoGrade 12

2 Answers

Ansh Mahapatra
ApprovedApproved Tutor Answer7 Years ago
Change in momentum=0.2*20=4 kg m/s
According to definition,
F=dp/dt
[Change in momentum/Time taken]
Therefore, F=4/0.5=8 N
Khimraj
ApprovedApproved Tutor Answer7 Years ago
Force = m\Deltav/\Deltat
SO F = 0.2*20/0.5 = 8N
Hope it clears.......................................................................................